# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 7.2 Dealing With Back Arcs
Example 7.2.
Figure 16: Illustration of the Decreasing Neighborhood Sequence Property (DNSP) where node 0 has more out-neighbors within rooted neighborhood 1 than out-neighbors in rooted neighborhood 2.
The definition of a back arc is an edge $(y,z)$, where $y\in R_{i}$, $z\in R_{j}$, and $j<i$, respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the rooted neighborhood organization, which contrasts with this.
Lemma 7.3 (No Back Arc) effectively outlines the fundamental problem with back arcs. 6.6 (DNSP Neighborhood Size Property) introduced how the DNSP impacted rooted neighborhoods, decreasing in size as we went further from the minimum out-degree node $v_{0}$. Then we were able to prove 6.8, which showed that as we moved further from the minimum out-degree node, these interior nodes were still keeping up their part of the load balance. Finally, we were able to combine these two lemmas with 6.4, which made each rooted neighborhood serve as its own SSNP and inside the $R_{i}$, we find that the interior degree would double. We can combine these three things, and we see that the back arcs lead to a node whose rooted neighborhood size is always greater than the head node of the back arc’s exterior neighborhood’s size.
@@
We assume here that all nodes in $G$ have the DSNP. Assume that $u_{i}\in R_{i}$ has its earliest back arc to $v_{k}\in R_{k}$ where $k<i$. A back arc $(u_{i},v_{k})$ means $v_{k}$ is a first out-neighbor of $u_{i}$. The neighbors of $v_{k}$ are then at distance two from $u_{i}$ through $v_{k}$ (a path of length two: $u_{i}\rightarrow v_{k}\rightarrow N^{+}(v_{k})$).
The back arc $(u_{i},v_{k})$ establishes a direct adjacency between $u_{i}$ and $v_{k}$, creating a path of length 1. However, the definition of $N^{++}(u_{i})$ requires paths of length 2 starting from $u_{i}$. The back arc does not create any directed paths of length 2 from $u_{i}$ to the out-neighbors of $v_{k}$. Therefore, the size of $|N^{++}(u_{i})|$ remains bounded below by $\text{deg}(v_{k})$. These second out-neighbors, which are the out-neighbors of $v_{k}$, which has a degree of $\text{deg}(v_{k})$, are in a neighborhood closer to $v_{0}$ than $u_{i}$. By ’first back arc,’ we mean the back arc originating from $u_{i}$ that targets the rooted neighborhood with the largest index $k$ such that $k<i$ and there are no back arcs from $u_{i}$ to any $R_{j}$ where $k<j<i$.
We also have that by the interior degree doubling of $u_{i}$ and the lack of back arcs before $i$, we know that $u_{i}$ will have only interior and exterior arcs. By $u_{i}$ having its interior degree double, if we can show that its exterior degree will double, then its overall degree will double. We just showed that $v_{k}$ was a second out-neighbor of $u_{i}$ through an exterior arc. By Lemma 6.6, we have that the number of nodes that $u_{i}$ can send through exterior arcs to $R_{i+1}$ is at most equal to the number of neighbors of $v_{k}$. That is, $|R_{i+1}|\leq|R_{k}|$.
@@
The proof above demonstrates that a node cannot have a back arc to an early neighborhood. To understand this proof, we must first investigate the assumptions. Our main assumption here is that the first back arc from $u_{i}$ is to the node $v_{k}$. As we saw in the proof, $v_{k}$ has more exterior out-neighbors than $u_{i}$, which is what causes $u_{i}$’s degree to double. We had to investigate the conditions surrounding $u_{i}$’s first and second out-neighbors and this back arc.
Secondly, we look at the notion of the first back arc. The neighborhoods are totally ordered. By mapping $u_{i}$ to its first back arc and showing that this back arc cannot exist, we have shown that $u_{i}$ cannot have a first back arc. Because $u_{i}$ cannot have a first back arc, this back arc of this type cannot exist.
xyu
Figure 17: Illustration of a transitive triangle with predecessor node $x$, successor nodes $y$ and $u$, where $u$ is also a successor of $y$
@@
7.3 Comparing GLOVER to BFS
When we contrast Breadth First Search (BFS) to the GLOVER data structure, we must understand that the former is a general purpose traversal algorithm whose goal is to traverse all arcs. The later is a domain-specific data structure that was designed specifically with the goal in mind of solving the Seymour Second Neighborhood Conjecture (SSNC).
One key difference between the two is Lemma 6.4 (Interior Degree Doubles), which GLOVER makes use of before beginning the traversal of any nodes. As noted, BFS traverses all kinds of arcs, whereas GLOVER does not traverse interior arcs because we already know that they double. The purpose of the traversal in the GLOVER algorithm is to traverse the exterior arcs in search of a node whose exterior also doubles. There are two possible cases, if such a node exists, we have a degree doubling node. If no such node exists, then we have proven the conjecture is false.
To illustrate the difference, consider Example 7.3, which demonstrates how back arcs can complicate SSNC analysis when using BFS.
Example 7.3.
xywz
Using GLOVER’s constraints, specifically Example 7.3 (No Back Arcs), prove that back arcs in oriented graphs will immediately produce degree-doubling nodes. This implies that the presence of the arc $(z,w)$, as shown in the diagram, which is a back arc, will indeed help solve the SSNC. The data structure is aided with the Decreasing Neighborhood Sequence Property, which helps to identify things like back arcs, whereas BFS would just see these as another edge in the graph.
While BFS provides a level-order traversal, GLOVER is specifically designed to enforce the constraints of oriented graphs and the SSNC. It actively seeks patterns that optimize and contradict these constraints, such as the back arc in the previous example. BFS will correctly label each node and its distance from the minimum out-degree node $(v_{0})$, without repeating any edge or node. The problem with this approach is that the BFS algorithm does not distinguish between valid paths—those adhering to oriented graph constraints—and paths created by back arcs. However, BFS treats all edges equally, whether they contribute to the level-ordered structure or create shortcuts via back arcs. This can obscure the violations of the GLOVER’s distance-based ordering that are crucial for identifying degree-doubling scenarios related to the SSNC.
In contrast, the GLOVER data structure was built for the purpose of the SSNC, towards proving the SSNC. The BFS algorithm, due to its general nature, is not designed to prove this conjecture. This distinction is crucial for both BFS and the GLOVER data structure, and ultimately, for the SSNC. BFS is a general traversal algorithm that was not designed to address the specific challenges of the SSNC.
Some of those patterns that emerge are, first, cycles within the rooted neighborhoods. For even-sized rooted neighborhoods, we are able to see symmetry about balance, uniformity, and a sense of order within each neighborhood. We must also make a point of the presence of cycles in each rooted neighborhood, which give rise to the interior degrees doubling, as well as the partitions about the minimum out-degree node that allow us to dissect a node’s degree into interior and exterior degree. These lemmas are all necessary components that take place before we could begin a traversal algorithm. This should distinguish the GLOVER setup, data structure and algorithm to come from BFS and similar procedures.
8 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our algorithm and show how it solves the SSNC, we need to address some potential problems and show that they will not hinder our algorithm.
8.1 Transitive Triangle
Here, we have $x\rightarrow y\rightarrow u$ and $x\rightarrow u$. The concept of a transitive triangle is important because it involves the risk of double-counting. The node $u$ is both a first and a second out-neighbor of $x$. When we improperly inflate a node’s neighborhood size with this double-counting, we risk calling $x$ a degree-doubling node even though its number of second out-neighbors is less than its number of first out-neighbors. This improper counting of second out-neighbors can lead to an erroneous algorithm.
In [2], Brantner et al., they showed that any graph without transitive triangles would necessarily contain a degree-doubling node. That node would be the minimum out-degree node $v_{0}$ in the graph $D$. By there being no transitive triangles in $D$, there would be no out-neighbors of this node to share the load of $v_{0}$ and stop it from having its degree double. We should note that in [2], it was not the intention to prevent the node $v_{0}$ from having its degree double. Instead, the absence of transitive triangles in $D$ provided them with the framework to find a degree-doubling node. While Brantner et al. used the absence of transitive triangles as a condition to guarantee a degree-doubling node, GLOVER leverages their presence in a structured way to facilitate load balancing and prevent degree doubling at earlier levels.
We contrast that with the GLOVER data structure, where the structure is built rooted neighborhood by rooted neighborhood on top of transitive triangles. Similar to the proof in Brantner et al., [2], we start with a minimum out-degree node, but the transitive triangles have a purpose - to prevent their parents and their ancestors from having their degrees double. This is combined with the Decreasing Neighborhood Sequence Property (DNSP) that forces the size of rooted neighborhoods to decrease as we move further from the minimum out-degree node $v_{0}$. This collapse in the cardinality of the rooted neighborhoods is matched by nodes that must meet requirements of having an interior out-degree of at least distance from the minimum out-degree node $v_{0}$. By Lemma 6.8 (General Load Balancing), if the node $u$ is in the rooted neighborhood $R_{i}$, then $u$ must have an interior degree at least $i$. Not only are these rooted neighborhoods getting denser as we move further out, but the number of transitive triangles is also increasing to support the parent nodes. The transitive arc $(x,u)$ means that $x$ has a direct connection to $u$, potentially reducing the ’load’ or the necessity for $y$ to be solely responsible for reaching $u$ and subsequent nodes. This sharing of responsibility helps prevent $x$’s out-degree from doubling.
The fact that the rooted neighborhoods have a finite amount of space leaves a limited amount of room for expansion from the minimum out-degree node. These transitive triangles, then, are not a hindrance, but a necessity that allows the GLOVER structure to continually perform in a predictable manner. This eventual termination of the decreasing sequence of positive rooted neighborhood sizes is where the load-balancing effect of transitive triangles becomes so apparent. It demonstrates how they partner with the DNSP to force these rooted neighborhoods to decrease. As the DNSP forces rooted neighborhoods to decrease in size, the increasing density (due to higher interior degrees and more transitive triangles) becomes crucial for maintaining connectivity and load balancing within the shrinking space, ultimately leading to a point where further expansion is impossible.
We were also able to partition the transitive triangles Brantner et. al. [2] into three distinct sets based on their characteristics in our GLOVER data structure. These sets include: those consisting of one parent and two children (which we term interior neighbors), those consisting of two parents and one child (resembling exterior neighbors), and those consisting of a parent, a child and a back arc (referred to as back arc triangles). Moreover, we established specific plans of action for addressing each of these transitive triangles types. The interior neighbors do not present a problem, as they inherently exhibit degrees-doubling properties. Exterior neighbors will be managed through a dedicated traversal algorithm. Furthermore, the presence of back arcs will immediately lead to the identification of a degree-doubling node, thus concluding our overall proof.
As a result, these precise definitions of interior neighbors, exterior neighbors and back arcs are not mere restatements of general transitive triangles. Instead, this partitioning provides the framework that transforms the graph theoretic challenge into actionable components for our algorithmic approach.
8.2 Seymour Diamond
The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is a four-vertex-oriented graph. What this situation presents is redundant second out-neighbors. There is a path $x\rightarrow y\rightarrow u$ in this diamond, and there is a path $x\rightarrow w\rightarrow u$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture. If we were simply counting nodes reachable in two steps from $x$, $u$ might be counted twice because it’s reached via both $y$ and $w$.
Definition 8.1.
Lemma 8.1.
Seymour Diamonds from Neighborhoods Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Neighborhood Sequence Property (DNSP). For any three consecutive rooted neighborhoods $R_{i}$, $R_{i+1}$, and $R_{i+2}$, let:
$u\in R_{i},\quad v,w\in R_{i+1},\quad y\in R_{i+2}$
The following arcs exist:
1. There are arcs $x\to y$ and $y\to u$,
2. There are arcs $x\to w$ and $w\to u$,
3. No back arcs exist between neighborhoods.
Through the GLOVER data structure, we introduced exterior neighbors. This offers a way to disambiguate second out-neighbors. Consider a scenario before the GLOVER structure, where a parent node $x$ has two children, $y$ and $w$, both with out-neighbor $z$. Generally because $z$ was reachable from $x$ through two different paths, it would be counted twice.
Now consider our GLOVER structure and how it would handle this same scenario. The exterior neighbors act as trackers, letting us know not just who the second out-neighbor is, but where it came from. Here, $\text{ext}(x,y)=z$ and $\text{ext}(x,w)=z$ say that we have two different paths to $z$. It also lets us know that we have one second out-neighbor $z$. The Seymour diamonds still exist in GLOVER, but these paths can be tracked and are not a deterrent. Not only is this not a problem, it’s one of the benefits of the GLOVER data structure. The exterior neighbor sets $\text{ext}(x,y)={z}$ and $\text{ext}(x,w)={z}$ indicate that while there are two paths to $z$, it is only a single unique second out-neighbor of $x$. Thus, GLOVER avoids over-counting.
xywu
Figure 18: Illustration of a Seymour diamond with predecessor node $x$, successor nodes $y$ and $w$, and successor of the successor of node $u$
Now that we have a better understanding of transitive triangles and Seymour diamonds and how they interact with our data structure. We can see that they are not a problem for the development of our algorithm. We will continue our investigation of oriented graphs. Next, we will combine all of the information into our primary findings. As we can see now, transitive triangles are essential components of the GLOVER’s design rather than barriers. Seymour diamonds, while not as essential as transitive triangles, are still not a stumbling block. The presence of these subgraphs represent the data structure’s underlying organization and has a direct impact on the algorithm’s functionality. By keeping track of the origins of second out-neighbors while using the concept of exterior sets first defined in this paper, GLOVER automatically eliminates the redundancies introduced by Seymour diamonds. This leverages the load-sharing capabilities of transitive triangles, transforming potential obstacles into integral components of its SSNC analytical framework.
@@ 9.1 Introduction
So far, this battle has been a battle between a catalyst, $v_{0}$, and the bottleneck, the Decreasing Neighborhood Sequence Property (DNSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each rooted neighborhood, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
@@ 9.2 Theorem Statement
Let $G$ be an oriented graph, with all nodes having the Decreasing Neighborhood Sequence Property (DNSP). Let $v_{0}$ be a minimum out-degree node where we set up our partition. Then Seymour’s conjecture holds true.
@@
By Lemma 6.6 (DNSP Neighborhood Size Property), the rooted neighborhoods $R_{i}$ strictly decrease in size. Since the sizes are positive integers, this sequence must eventually reach a point where $|R_{i}|\leq 2$. These are special cases because load sharing can no longer take place among the children.
Case 1: If $|R_{i}|=1$, let $R_{i}={v}$. Then $u_{i-1}$ has an exterior out-neighbor $v$. Since $|R_{i}|$ is minimal, $|R_{i-1}|>1$. By Lemma 6.8 (General Load Balancing), $v$ should have an interior degree of at least $i$. However, since $|R_{i}|=1$, this is impossible. This implies a failure of the DNSP at some earlier node, leading to degree doubling.
Case 2: If $|R_{i}|=2$, say $R_{i}={v_{1},v_{2}}$, then any predecessor $u_{i-1}$ had at least two exterior neighbors. For $v_{1}$ and $v_{2}$ to maintain the load balancing, they would need interior connections. However, with only two nodes, at least one node must have all its outgoing arcs to $R_{i+1}$ or back to an earlier neighborhood, thus forcing a degree-doubling scenario for its predecessor in $R_{i-1}$. ∎
@@
What we see from this table is that once the size of $R_{2}$ reached only 2 nodes, we were guaranteed that some node in $R_{1}$ would have its degree dou
@@
Degree-Doubling by Rooted Neighborhood Density In an oriented graph $G$ with a minimum out-degree node $v_{0}$ and rooted neighborhoods $R_{1},\allowbreak R_{2},\allowbreak\ldots\allowbreak R_{n}$, as the neighborhoods shrink in size as their distance from the minimum out-degree node increases, and as the required interior degrees rise, a collision is inevitable.
There are two possibilities:
  • Case 1: The rooted neighborhoods continue to shrink in size. Simultaneously, the interior degrees continue to grow. We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
  • Case 2: For small minimum out-degree nodes, the rooted neighborhood’s shrinkage is not fast enough to cause this collision. For instance, if the minimum degree $\delta$ is very small, the initial rooted neighborhoods might be small, and the linear increase in required interior degree might not create a collision before the graph runs out of nodes.
@@
The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one of abundance. There were many who doubted that one even existed. Now there are many.
@@ 9.3 Algorithm
Before we formally present the algorithm we have described in this paper, we will give two additional lemmas that will help with both the algorithm and the complexity of the algorithm. This algorithm leverages the GLOVER data structure and the properties derived from the assumption of the Decreasing Neighborhood Sequence Property (DNSP) to detect a degree-doubling node. It systematically checks for back arcs, violations of the interior degree requirements, and the eventual shrinkage of rooted neighborhoods to a size that forces degree doubling in a predecessor.
@@
One Interior Fail Means All Fail Assume that for a node $u_{i}\in R_{i}$, the condition $|\text{int}(u_{p1},u_{i})|\geq i$ fails for some $u_{p1}\in R_{i-1}$. Then for all other $u_{p2}\in R_{i-1}$, the same condition $|\text{int}(u_{p2},u_{i})|\geq i$ fails, making every node in $R_{i-1}$ a degree-doubling node.
@@
Assume, for contradiction, that there exists a node $u_{i}\in R_{i}$ and a parent $u_{p1}\in R_{i-1}$ such that $|\text{int}(u_{p1},u_{i})|<i)$.
This implies that $u_{i}$ sends more than $\delta-i$ arcs to nodes in the next neighborhood $R_{i+1}$, i.e., $|\text{ext}(u_{i},R_{i+1})|\geq\delta-i+1$.
Now let $u_{p2}\in R_{i-1}$ be any other parent of $u_{i}$. Since arcs are oriented and back arcs are excluded by construction, the neighbors of $u_{i}$in $R_{i+1}$ must be second out-neighbors of $u_{p2}$. Thus, $|\text{ext}(u_{i},R_{i+1})|\leq|N^{++}(u_{p2})|$.
But since $|\text{ext}(u_{i},R_{i+1})|\geq\delta-i+1$, and since $u,p2$ has at most $\delta-i$ first out-neighbors (i.e., in $R_{i}$), we get:
$|N^{++}(u_{p2})|\geq\delta-i+1>\delta-i\geq|N^{+}(u_{p2})|.$
Therefore, $|N^{++}(u_{p2})|\geq|N^{+}(u_{p2})|,$ which means $u_{p2}$ is a degree-doubling node, violating the DNSP.
Since this argument holds for any other parent $u_{p2}$ of $u_{i}$, all such parents become degree-doubling nodes. Thus, one interior failure at $u_{i}$ implies that all its parents in $R_{i-1}$ fail the interior degree condition.
@@
One Interior Succeed Means All Succeed Assume that for a node $u_{p1}\in R_{i-1}$ and all nodes $u_{i}\in R_{i}$ we have $|\text{int}(u_{p1},u_{i})|\geq i$. Then for all other $u_{p2}\in R_{i-1}$ the same condition $|\text{int}(u_{p2},u_{i})|\geq i$ holds, and no node in $R_{i-1}$ is a degree-doubling node.
@@
Assume that for a particular parent node $u_{p1}\in R_{i-1}$, the interior arc condition holds for every child node $u_{i}inR_{i}$:
$|\text{int}|(u_{p1},u_{i})|\geq i.$
By definition, this implies that each such node $u_{i}$ sends at most $\delta-i$ arcs to the next neighborhood $R_{i+1}$, i.e., $|\text{ext}|(u_{i},R_{i+1})|<\delta-i$.
Now, consider any other parent node $u_{p2}\in R_{i-1}$. We want to show that for every shared child $u_{i}$, the interior condition also holds for $u_{p2}$.
Note that the arcs from $u_{i}$ to $R_{i+1}$ are the only possible contributors to $N^{++}(u_{p2})$, since $u_{p2}$ has no direct arcs to $R_{i+1}$. That is:
$N^{++}(u_{p2})\subseteq\bigcup_{u_{i}\in N^{+}(u_{p}2)}\text{ext}|(u_{i},R_{i+1})$
Because $|\text{ext}|(u_{i},R_{i+1}|<\delta-i$, and the total number of nodes in $R_{i+1}$ is also bounded by $|R_{i+1}|<\delta-i$ (by Lemma 6.7), it follows that $N^{++}(u_{p2})$ cannot exceed $N^{+}(u_{p2})$ in cardinality.
Thus,
$|N^{++}(u_{p2})|<|N^{+}(u_{p2})|,$
meaning $u_{p2}$ is not a degree-doubling node.
Working backward, this implies that the number of arcs each $u_{i}$ sends to $R_{i+1}$ must also be bounded by $\delta-i$, just as for $u_{p1}$. Therefore, the interior arc condition must hold for all parents of $u_{i}$, including $u_{p2}$. ∎
@@
1:Determine a minimum out-degree node $v_{0}\in V$.
2:Partition the set of vertices $V$ into ordered sets $R_{0},R_{1},\ldots,R_{k}$ based on their out-distance from $v_{0}$, where $R_{0}=\{v_{0}\}$ and $R_{i}=N^{+}(R_{i-1})\setminus\bigcup_{j=0}^{i-1}R_{j}$ for $i>0$.
@@
5:         if there exists an arc $(u_{i},w)\in E$ such that $w\in R_{k}$ where $k<i$ then
[…]
@@
10:                  if $|\text{int}(u_{i-1},u_{i})|<i$ then
[…]
12:                       HALT
13:                  end if
@@
16:     end for
17:     if $|R_{i}|\leq 2$ then
18:         if $i>0$ then
[…]

### v12

@@ 7.2 Dealing With Back Arcs
The definition of a back arc is an edge $y\to z$, where $y\in R_{i}$, $z\in R_{j}$, and $j<i$, respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the rooted neighborhood organization, which contrasts with this.
Lemma 7.3 (No Back Arc) effectively outlines the fundamental problem with back arcs. Lemma 6.8 (DNSP Impact on Neighborhood Size) introduced how the DNSP impacted rooted neighborhoods, decreasing in size as we went further from the minimum out-degree node $v_{0}$. Then we were able to prove Lemma 6.7 (Generalized Load Balance), which showed that as we moved further from the minimum out-degree node, these interior nodes were still keeping up their part of the load balance. Finally, we were able to combine these two lemmas with Lemma 6.4 (Interior Degree Doubling), which made each rooted neighborhood serve as its own SSNP and inside the $R_{i}$, we find that the interior degree would double. When we combine these three lemmas, we see that the back arcs lead to a node whose rooted neighborhood size is always greater than the head node of the back arc’s exterior neighborhood’s size.
@@
We assume here that all nodes in $G$ have the DNSP. Assume that $u_{i}\in R_{i}$ has its earliest back arc to $v_{k}\in R_{k}$ where $k<i$. A back arc $u_{i}\to v_{k}$ means $v_{k}$ is a first out-neighbor of $u_{i}$. The neighbors of $v_{k}$ are then at distance two from $u_{i}$ through $v_{k}$ (a path of length two: $u_{i}\to v_{k}\to N^{+}(v_{k})$).
By ’first back arc,’ we mean the back arc originating from $u_{i}$ that targets the rooted neighborhood with the largest index $k$ such that $k<i$ and there are no back arcs from $u_{i}$ to any $R_{j}$ where $k<j<i$.
We also have that by the interior degree doubling of $u_{i}$ and the lack of back arcs before $i$, we know that $u_{i}$ will have only interior and exterior arcs. By $u_{i}$ having its interior degree double, if we can show that its exterior degree will double, then its overall degree will double. We just showed that $w\in N^{+}(v_{k})$ was a second out-neighbor of $u_{i}$ through an exterior arc. By Lemma 6.9 (DNSP Impact on Neighborhood Size), we have that $\text{ext}(u_{i-1},u_{i})<\delta-i$, for some parent $u_{i-1}\in R_{i-1}$ of $u_{i}$. Similarly, because $v_{0}$ is a minimum out-degree node in $G$ we have that $d^{+}(v_{k})=|N^{+}(v_{k})|\geq\delta$. We can combine these two inequalities and we see that
$|N^{+}(v_{k})|\geq\delta>\delta-i>\text{ext}(u_{i-1},u_{i}).$
Simplifying, we see that
$|N^{+}(v_{k})|>\text{ext}(u_{i-1},u_{i}).$
Even the situation where $v_{k}\in R_{i-1}$ would not stop $u_{i}$ from becoming a degree-doubling node because Lemma 6.7 (Generalized Load Balance) states that $i$ of $u_{i}$’s neighbors must be interior neighbors. This causes the inequality to hold.
@@
The proof above demonstrates that a back arc to an early neighborhood would simply lead to a degree-doubling node. Our main assumption here is that the first back arc from $u_{i}$ is to the node $v_{k}$. What is essential is the Graph Level Order that allows us to organize, not only the nodes themselves, but also the back arcs in a schematic way so as to be able to select the earliest one from $u_{i}$. Similarly, this data structure has been allowing us to define arcs by their relationship of their nodes to the distances of their rooted neighborhoods. Nodes with the same distances and the same parent are interior neighbors, nodes with a difference of 1 and the same parent are exterior neighbors, and nodes with a difference greater than 1 and the same parent are called back arcs. What the proof shows is that $v_{k}$ has a higher degree than $u_{i}$’s number of exterior neighbors, which is what causes $u_{i}$’s degree to double.
@@ 8.1 Introduction
So far, this battle has been a battle between a catalyst, $v_{0}$, and the bottleneck, the Decreasing Neighborhood Sequence Property (DNSP). We have proved that: Back arcs cannot exist; interior degrees must double inside each rooted neighborhood, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
@@ 8.2 Theorem Statement
Let $G$ be an oriented graph where all nodes satisfy the Decreasing Neighborhood Sequence Property (DNSP). Let $v_{0}$ be a minimum out-degree node used to set up our rooted neighborhood partition. Then Seymour’s conjecture holds true for $G$.
@@
By Lemma 6.8 (DNSP Neighborhood Size Property), the sizes of rooted neighborhoods $|R_{1}|,|R_{2}|,\ldots$ form a strictly decreasing sequence of positive integers. Consequently, there exists a smallest $i>0$ such that
$|R_{i}|\leq 2.$
Let $u_{i-1}\in R_{i-1}$ be a parent of the nodes in $R_{i}$.
Case 1: If $|R_{i}|=1$, say $R_{i}=\{v\}$.
Then $u_{i-1}$ has an exterior out-neighbor $v$. Since $|R_{i}|$ is minimal, $|R_{i-1}|>1$. By Lemma 6.7 (General Load Balancing), $v$ should have an interior degree of at least $i$. However, with only one node in $R_{i}$, this is impossible without violating the DNSP at some previous node, which leads to degree doubling.
Case 2: If $|R_{i}|=2$, say $R_{i}=\{v_{1},v_{2}\}$.
Then the parent $u_{i-1}$ had at least two exterior neighbors. For $v_{1}$ and $v_{2}$ to maintain the load balancing, they would need interior connections. However, with only two nodes, at least one node must have all its outgoing arcs to $R_{i+1}$ or back to earlier neighborhoods, forcing a degree-doubling scenario for its predecessor in $u_{i-1}$. ∎
@@
What we see from this table is that once the size of $R_{2}$ reached only 2 nodes, we were guaranteed that some node in $R_{1}$ would have its degree double by the combination of size of $R_{2}$ and the interior degrees doubling. The nodes in $R_{1}$ have interior degrees of 1 and thus automatically have their degrees double in that neighborhood. The traversal algorithm would simply search for a node whose degree is less than 2. This is achieved in $R_{2}$ where at least one node must have the exterior degree of 3. This implies that all nodes in $R_{1}$ that are parents of this node will have their degrees double.
@@
Degree-Doubling by Rooted Neighborhood Density In an oriented graph $G$ with a minimum out-degree node $v_{0}$ and rooted neighborhoods $R_{1},\allowbreak R_{2},\allowbreak\ldots\allowbreak R_{n}$, the shrinkage of neighborhoods and growth of interior degrees guarantees an inevitable collision.
Two scenarios emerge:
  • Case 1: The rooted neighborhoods shrink rapidly while interior degrees grow.
    Eventually, no oriented graph can support the required structure, forcing degree doubling.
  • Case 2: For very small minimum degree $\delta$, the neighborhoods shrink slowly and interior degree growth does not immediately force a collision before nodes run out.
@@
The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one of abundance. There were some who doubted that one even existed. Now there are many.
@@ 8.3 Algorithm
Before we formally present the algorithm we have described in this paper, we will give two additional lemmas that will help with both the algorithm and the complexity of the algorithm. This algorithm leverages the Graph Level Order and the properties derived from the assumption of the Decreasing Neighborhood Sequence Property (DNSP) to detect a degree-doubling node. It systematically checks for back arcs, violations of the interior degree requirements, and the eventual shrinkage of rooted neighborhoods to a size that forces degree doubling in a predecessor.
@@
One Interior Fail Means All Fail
Let $u_{i}\in R_{i}$ be a node such that for some parent $u_{p_{1}}\in R_{i-1}$, the interior degree condition fails:
$|\text{int}(u_{p1},u_{i})|<i.$
Then the same condition fails for every other parent $u_{p_{2}}\in R_{i-1}$, implying that every node in $R_{i-1}$ is a degree-doubling node.
@@
Suppose, for contradiction, there exists a node $u_{i}\in R_{i}$ and a parent $u_{p_{1}}\in R_{i-1}$ with
$|\text{int}(u_{p_{1}},u_{i})|<i).$
This implies that $u_{i}$ sends more than $\delta-i$ arcs to nodes in the next neighborhood $R_{i+1}$, i.e.,
$|\text{ext}(u_{i},R_{i+1})|\geq\delta-i+1.$
Consider any other parent $u_{p_{2}}\in R_{i-1}$ of $u_{i}$. Since arcs are oriented and back arcs are excluded, the neighbors of $u_{i}$ in $R_{i+1}$ are second out-neighbors of $u_{p2}$. Thus,
$|\text{ext}(u_{i},R_{i+1})|\leq|N^{++}(u_{p_{2}})|.$
But since
$|\text{ext}(u_{i},R_{i+1})|\geq\delta-i+1,$
and since $u_{p_{2}}$ has at most $\delta-i$ first out-neighbors (in $R_{i}$), we have
$|N^{++}(u_{p_{2}})|\geq\delta-i+1>\delta-i\geq|N^{+}(u_{p_{2}})|.$
Therefore,
$|N^{++}(u_{p_{2}})|\geq|N^{+}(u_{p_{2}})|,$
which means $u_{p_{2}}$ is a degree-doubling node, violating the DNSP.
Since this argument holds for $u_{p_{2}}\in R_{i-1}$, all parents of $u_{i}$ fail the interior degree condition, and thus all become degree-doubling nodes.
@@
One Interior Succeed Means All Succeed Suppose that for a node $u_{p_{1}}\in R_{i-1}$ and every $u_{i}\in R_{i}$ the interior degree condition holds:
$|\text{int}(u_{p_{1}},u_{i})|\geq i.$
Then this condition holds for all other parents $u_{p_{2}}\in R_{i-1}$, and no node in $R_{i-1}$ is degree doubling.
@@
Assume that for a particular parent node $u_{p_{1}}\in R_{i-1}$, the interior arc condition holds for every child node $u_{i}\in R_{i}$:
$|\text{int}(u_{p1},u_{i})|\geq i.$
By definition, this implies that each such node $u_{i}$ sends at most $\delta-i$ arcs to the next neighborhood $R_{i+1}$, i.e.,
$|\text{ext}(u_{i},R_{i+1})|<\delta-i.$
Now, for any other parent $u_{p_{2}}\in R_{i-1}$, We want to show that for every shared child $u_{i}$, the interior condition also holds for $u_{p_{2}}$.
$N^{++}(u_{p_{2}})\subseteq\bigcup_{u_{i}\in N^{+}(u_{p_{2}})}\text{ext}(u_{i},R_{i+1}).$
Since each $u_{i}$ sends at most $\delta-i$ arcs to $R_{i+1}$, and by Lemma 6.9, the size of $R_{i+1}$ is also bounded by $\delta-i$, it follows that
$|N^{++}(u_{p2})|\leq|N^{+}(u_{p_{2}})|\times(\delta-i)\leq|N^{+}(u_{p_{2}}|\times 1,$
where the factor 1 applies if the out-degree matches the bound tightly. More precisely, the cardinality of second out-neighbors does not exceed that of first out-neighbors.
Hence,
$|N^{++}(u_{p_{2}})|<|N^{+}(u_{p_{2}})|,$
and $u_{p_{2}}$ is not a degree-doubling node.
Working backward, since this holds for all $u_{p_{2}}\in R_{i-1}$, the interior degree condition
$|\text{int}(u_{p_{2}},u_{i})|\geq i.$
must hold for all parents $u_{p_{2}}$ and children $u_{i}$. ∎
@@
1:Determine a minimum out-degree node $v_{0}\in G$.
2:Partition $V$ into ordered sets $R_{0},R_{1},\dots,R_{k}$ where
$R_{0}=\{v_{0}\},\quad R_{i}=N^{+}(R_{i-1})\setminus\bigcup_{j=0}^{i-1}R_{j}\quad\text{for }i>0.$
@@
5:         if there exists an arc $u_{i}\to w\in G$ with $w\in R_{m}$ where $m<i$ then
6:              Mark $u_{i}$ as degree_doub[back]
@@
10:                  Let $u_{p}\in R_{i-1}$ be a representative parent of $u_{i}$.
11:                  if $|\text{int}(u_{p},u_{i})|<i$ then
12:                       Mark $u_{p}$ as degree_doub[dense]
13:                       HALT
@@
16:         end if
17:     end for
18:     if $i>0$ and $|R_{i}|\leq 2$ then
19:         Mark $u_{p}\in R_{i-1}$ as degree_doub[size]